### 05 Dec 2023 v1

# Friedmann equations in the Codazzi parametrization of Cotton and extended theories of gravity and the Dark Sector

Abstract
The Friedmann equations of Cotton gravity provide a simple parametrization to reproduce, by tuning a single function, the Friedmann equations of several extensions of gravity, such as $f(R)$, modified Gauss-Bonnet $f(G)$, teleparallel $f(T)$, and more. It also includes the recently proposed Conformal Killing gravity in FRW space-times. The extensions generally have the form of a Codazzi tensor that may be associated to the dark sector. Fixing it by a suitable equation of state accomodates most of the postulated models that extend $\Lambda$CDM, as the Chevallier-Polarski-Lindler model.
Cotton gravity; Extended theories of gravity; Codazzi tensor; Conformal Killing gravity; Friedmann-Robertson-Walker spacetime
I Introduction
In recent years there has been a flourishing of extended theories of gravity to address the problem of the Dark Sector. They modify the Einstein equations by adding a term $H_{jk}$ to the energy-momentum tensor $T_{jk}$ of matter:
$R_{jk}-\frac{1}{2}g_{jk}R=\kappa(T_{jk}+H_{jk})$ (1)
The term originates from a new form of gravitational action or new particles.
Large-scale cosmology is staged in Friedmann-Robertson-Walker (FRW) space-times, where the Weyl tensor $C_{jklm}$ is zero. This fact ushers Codazzi tensors.
$\bullet$ If $\nabla_{m}C_{jkl}{}^{m}=0$, then
$R_{ij}-\frac{R}{2}g_{ij}=S_{ij}-g_{ij}S^{k}{}_{k}$ (2)
where the Schouten tensor $S_{ij}=R_{ij}-\frac{1}{6}Rg_{ij}$ is a Codazzi tensor, i.e. $\nabla_{i}S_{jk}=\nabla_{j}S_{ik}$.
This means that $T_{jk}+H_{jk}$ has the same decomposition. The Codazzi condition ensures that $\nabla^{k}(T_{kl}+H_{kl})=0$.
Moreover, in a FRW space-time the sum must have the perfect fluid structure of the Einstein tensor.
The vast majority of extended models of gravity in FRW space-times specify this property for the radiation-matter sector, with conservation. This entails a Codazzi decomposition of the perfect fluid tensor. Then, necessarily, despite the often complex stucture of the tensor $H_{kl}$, the dark sector is perfect fluid and conserved.
For these models:
$H_{jk}=\mathscr{C}_{jk}-g_{jk}\mathscr{C}^{p}{}_{p},$ (3)
$\nabla_{i}\mathscr{C}_{jk}=\nabla_{j}\mathscr{C}_{ik}$ (4)
The aim of this work is to uncover this common structure, albeit the different origins of the various cosmological models. We explicitly show this in plenty of well studied extended gravity models in FRW space-times: $f(R)$, Gauss-Bonnet $f(G)$, teleparallel $f(T)$, Einsteinian cubic $f(P)$, Conformal Killing gravity (CKT), Lovelock.
An inclusive and simple model which they fit in, stems from Cotton gravity.
In 2021 Junpei Harada [20] introduced a modification of General Relativity (GR) named “Cotton gravity” (CG), with field equations
$C_{jkl}=\nabla_{j}T_{kl}-\nabla_{k}T_{jl}-\frac{1}{3}\left(g_{kl}\nabla_{j}T-g_{jl}\nabla_{k}T\right)$ (5)
$T_{kl}$ is the matter energy-momentum tensor with trace $T$ and $C_{jkl}$ is the Cotton tensor:
$C_{jkl}\equiv\nabla_{j}\left(R_{kl}-\frac{R}{6}g_{kl}\right)-\nabla_{k}\left(R_{jl}-\frac{R}{6}g_{jl}\right)$ (6)
It is $C_{jkl}=-2\nabla_{m}C_{jkl}{}^{m}$. The property $g^{kl}C_{jkl}=0$ implies that $\nabla^{p}T_{jp}=0$. Cotton gravity was devised so that any solution of GR is a solution of CG.
Soon after, Harada applied his theory to describe the rotation curves of 84 galaxies without assuming the presence of dark matter [21]. The general spherically symmetric static vacuum solution was then obtained by Gogberashvili and Girvliani [18], with a long range modification of Newton’s law. A static solution of Cotton gravity with electric and magnetic charges was obtained in ([32], eq.80)
An important progress was made in [31] in showing that the equations of Cotton gravity are equivalent to the standard GR equations corrected by an arbitrary Codazzi tensor
$R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+\mathscr{C}_{kl}-g_{kl}\mathscr{C}^{r}{}_{r}$ (7)
$\nabla_{j}\mathscr{C}_{kl}=\nabla_{k}\mathscr{C}_{jl}$
This is precisely the statement in eq.(3), in the frame of FRW solutions. Cotton gravity exhibits the maximal freedom in specifying $H_{kl}$. We refer to eq.(7) as the “Codazzi parametrization”.
While Harada’s equations have third order derivatives of the metric, the equivalent equations (7) are second order.
Sussman and Nájera [43] used (7) to produce FRW solutions of Cotton gravity. They posed a modified Friedmann equation with a scalar function $\mathcal{K}(t)$ and obtained the components of a perfect fluid Codazzi tensor for which eq.(7) are satisfied. Very recently, they published a paper [44] with several non-trivial CG solutions that generalise the well known GR solutions: FLRW, Lemaitre-Tolman-Bondi and Szekeres, as well as static perfect fluid spherically symmetric solutions (with application to galactic rotation curves) and and non-static shearfree.
Also motivated by this result, we propose a general discussion based upon Theorem 2.1 in [31]: a FRW space-time always contains a perfect fluid Codazzi tensor that displays a freedom in its parameters.
In Section 2 we write the Friedmann equations for Cotton gravity in a FRW space-time.
In Section 3 we recognize that some of the most important extended theories of gravity have the following intriguing property: their Friedmann equations coincide with those of Cotton gravity by a suitable choice of the Codazzi tensor. We show this explicitly by providing the specific Codazzi tensor for $f(R)$ gravity, Gauss-Bonnet $f(G)$ gravity, $f(T)$ gravity, cubic Einsteinian and $f(P)$ gravity, Lovelock.
In Section 4 we show that the Codazzi parametrization of CG extends the recently introduced conformal Killing gravity [22, 23, 33], at least in FRW space-times.
In Section 5 the dark sector is fixed by requesting an EoS. It accomodates in a unified description the best known redshift dependent models, as the Chevallier-Polarski-Lindler model.
Notation: $i,j,k,\ldots=0,1,2,3$, $\mu,\nu,\ldots=1,2,3$. A dot operator $\dot{X}=u^{k}\nabla_{k}X$ is the time derivative in the comoving frame defined by $u^{0}=1$, $u^{\mu}=0$. $X_{[ijk]}$ is the cyclic sum $X_{ijk}+X_{kij}+X_{jki}$.
II Friedmann equations of Cotton gravity in FRW space-times
Generalized Robertson Walker space-times (GRW) are Lorentzian manifolds that extend FRW space-times with metric
$ds^{2}=-dt^{2}+a(t)^{2}g_{\mu\nu}^{\star}({\bf x})dx^{\mu}dx^{\nu},$ (8)
where $g_{\mu\nu}^{\star}({\bf x})$ is a positive definite metric and $a(t)$ is the scale factor. A covariant characterization is the existence of a vector field $u_{k}u^{k}=-1$ that is shear-free, vorticity-free and acceleration-free, and eigenvector of the Ricci tensor [30], i.e.
$\nabla_{j}u_{k}=H(g_{jk}+u_{j}u_{k}),$ (9)
$R_{ij}u^{j}=\xi u_{i}$ (10)
where $H=\dot{a}/a$ is Hubble’s parameter, $\xi=3(H^{2}+\dot{H})=3\ddot{a}/a$. The condition (10) is equivalent to $\nabla_{j}H=-\dot{H}u_{j}$. Its divergence and the contracted Bianchi identity give
$\dot{R}-2\dot{\xi}=-2H(R-4\xi)$ (11)
whose solution is [11]
$R=\frac{R^{\star}}{a^{2}}+12H^{2}+6\dot{H}$ (12)
where $R^{\star}$ is the spatial curvature. In $d=4$ and whenever $C_{jklm}u^{m}=0$ the GRW spacetime is a FRW space-time.
In a FRW space-time the natural form of the Codazzi tensor in eq.(7) is perfect fluid. $\Lambda g_{kl}$ with a constant $\Lambda$, is trivially a Codazzi tensor.
This simple result is proven in [31] (theorem 2.1):
Proposition 1
In a FRW space-time the tensor
$\mathscr{C}_{kl}=\mathscr{A}u_{k}u_{l}+\mathscr{B}g_{kl}+\frac{\Lambda}{3}g_{kl}$ (13)
is Codazzi provided that: $\nabla_{j}\mathscr{A}=-\dot{\mathscr{A}}u_{j}$, $\nabla_{j}\mathscr{B}=-\dot{\mathscr{B}}u_{j}$,
˙$\dot{\mathscr{B}}=-H\mathscr{A}$ (14)
Proof. The first two conditions mean that $\mathscr{A}=\mathscr{A}(t)$ and $\mathscr{B}=\mathscr{B}(t)$. Eq.(14) requires ˙$\dot{\mathscr{B}}\neq 0$. Next, with $\nabla_{j}\mathscr{A}=-\dot{\mathscr{A}}u_{j}$, $\nabla_{j}\mathscr{B}=-\dot{\mathscr{B}}u_{j}$, eqs.(9) and (14) it is:
$\nabla_{j}\mathscr{C}_{kl}=-u_{j}u_{k}u_{l}(\dot{\mathscr{A}}+2\dot{\mathscr{B}})-\dot{\mathscr{B}}(u_{l}g_{jk}+u_{k}g_{jl}+u_{j}g_{kl})$
Therefore (13) is a Codazzi tensor for any choice of the scale factor. $\square$.
It implies that any FRW space-time is a solution of Cotton gravity with (13), and leaves an interesting degree of freedom $\mathscr{B}$ in choosing the Codazzi tensor.
Eq.(7) is written with the input (13), the general form of the Ricci tensor of a FRW space-time
$R_{kl}=\frac{1}{3}(R-4\xi)u_{l}u_{k}+\frac{1}{3}(R-\xi)g_{kl}$
and the stress energy tensor $T_{kl}=(\mu+p)u_{l}u_{k}+pg_{kl}$ with energy density $\mu$ and pressure $p$ of ordinary matter.
Contractions with $u^{k}u^{l}$ and $g^{kl}$ and a simple rearrangement provide the Friedmann equations of Cotton gravity in a FRW space-time
$\kappa\mu=\frac{R}{2}-\xi-3\mathscr{B}-\Lambda$ (15)
$\kappa p=-\frac{R}{6}-\frac{\xi}{3}+3\mathscr{B}+\frac{\dot{\mathscr{B}}}{H}+\Lambda$ (16)
They are the standard Friedmann equations of GR augmented by the Codazzi terms. Such terms naturally correspond to the Dark Sector:
$H_{kl}=$ $(\mathscr{A}u_{k}u_{l}+\mathscr{B}g_{kl})-g_{kl}(4\mathscr{B}-\mathscr{A})$ (17)
$\equiv$ $(\mu_{D}+p_{D})u_{k}u_{l}+g_{kl}p_{D}$
The function $\mathscr{B}(t)$ parametrizes the energy density and the pressure of the dark sector:
$\mu_{D}=3\mathscr{B}$ (18)
$p_{D}=-3\mathscr{B}-\dot{\mathscr{B}}/H$ (19)
It implies the conservation law $\dot{\mu}_{D}=-3H(p_{D}+\mu_{D})$ coming from $\nabla_{k}H^{k}{}_{j}=0$ or the equivalent Codazzi condition for $\mathscr{A}u_{k}u_{l}+\mathscr{B}g_{kl}$ in FRW spacetimes.
III Reproducing the Friedmann equations of extended theories
We show that the Friedmann equations (15) and (16) of Cotton gravity may reproduce the Friedmann equations of other extended theories in absence of cosmological constant. With eq.(12) and $\xi=3(H^{2}+\dot{H})$ we write them as
$\kappa\mu=\frac{R^{\star}}{2a^{2}}+3H^{2}-3\mathscr{B}$ (20)
$\kappa p=-\frac{R^{\star}}{6a^{2}}-3H^{2}-2\dot{H}+3\mathscr{B}+\frac{\dot{\mathscr{B}}}{H}$ (21)
The comparison with the Friedmann equations of other theories selects the function $\mathscr{B}(t)$ that reproduces them.
III.1 $\mathbf{f(R)}$ gravity
Perhaps it is the best known extended theory of gravity. It was introduced by Buchdahl in 1970 [3] and gained popularity with the works on cosmic inflation by Starobinsky [41]. Recently $f(R)$ theories are possible candidates to explain the observed cosmic acceleration [8]. The action of $f(R)$ gravity is
$S=\frac{1}{2\kappa}\int d^{4}x\sqrt{-g}f(R)+S^{(m)}$
where $S^{(m)}$ is the matter term. With the notation $f_{R}=df/dR$, the field equations are [40]:
$f_{R}R_{kl}-\frac{f}{2}g_{kl}-[\nabla_{k}\nabla_{l}-g_{kl}\square]f_{R}=\kappa T_{kl}$ (22)
They can be rewritten in the form (1). In [9] it was proven that in a FRW space-time the resulting term $H_{jk}$ is a perfect fluid tensor.
For the spatially flat ($R^{\star}=0)$ FRW space-time the Friedmann equations of $f(R)$ gravity are ([40], eqs.75, 76):
$\kappa\mu=3f_{R}H^{2}-\tfrac{1}{2}(Rf_{R}-f)+3H\dot{R}f_{RR}$ (23)
$(3H^{2}+2\dot{H})f_{R}=-\big{[}\kappa p+\dot{R}^{2}f_{RRR}+2H\dot{R}f_{RR}$ (24)
$\qquad\qquad\qquad\qquad+\ddot{R}f_{RR}+\tfrac{1}{2}(f-Rf_{R})\big{]}$
In comparing (23) with (20) we identify
$\mathscr{B}=H^{2}(1-f_{R})+\tfrac{1}{6}(Rf_{R}-f)-H\dot{R}f_{RR}$ (25)
In computing $\dot{\mathscr{B}}$ we note that $\dot{f}(R)=f_{R}(R)\dot{R}$, $\dot{f_{R}}(R)=f_{RR}\dot{R}$, $\dot{f}_{RR}=f_{RRR}\dot{R}$, so that
˙$\dot{\mathscr{B}}=2H\dot{H}(1-f_{R})+f_{RR}(\frac{R}{6}-\dot{H}-H^{2})\dot{R}$
$-H\ddot{R}f_{RR}-H\dot{R}^{2}f_{RRR}$
The restriction $R^{\star}=0$ in (12) gives $R=12H^{2}+6\dot{H}$.
We obtain $-\mathscr{A}$:
˙$\frac{\dot{\mathscr{B}}}{H}=2\dot{H}(1-f_{R})+(H\dot{R}-\ddot{R})f_{RR}-\dot{R}^{2}f_{RRR}$ (26)
Using now (21) we obtain (24).
Proposition 2
The Friedmann equations of Cotton gravity with the perfect fluid Codazzi tensor (13) are the Friedmann equations of $f(R)$ gravity with the choice (25).
III.2 $\mathbf{f(G)}$ gravity
A second well known extended theory that tries to solve the problem of dark energy is the Gauss-Bonnet gravity, alias $f(G)$ gravity [37, 38, 13].
$S=\int d^{4}x\sqrt{-g}\left[\frac{R}{2\kappa}+f(G)\right]+S^{(m)}$
where $G=R^{2}-4R_{kl}R^{kl}+R_{jklm}R^{jklm}$ is the Gauss-Bonnet invariant.
The field equations may be written in the form $R_{kl}-\frac{1}{2}Rg_{kl}=\kappa(T_{kl}+H_{kl})$ with the following divergence-free tensor $H_{kl}$ [10, 19]:
$H_{kl}$ $=\tfrac{1}{2}g_{kl}f-2f_{G}(RR_{kl}-4R_{kq}R_{l}^{q}+2R_{k}\,^{pqr}R_{lpqr})$
$-4f_{G}R_{k}\,^{pq}\,{}_{l}R_{pq}+2R(\nabla_{k}\nabla_{l}f_{G}-g_{kl}\square f_{G})$
$-4(R_{l}{}^{p}\nabla_{p}\nabla_{k}f_{G}+R_{k}{}^{p}\nabla_{p}\nabla_{l}f_{G})+4(\square f_{G})R_{kl}$
$+4(R^{pq}g_{kl}-R_{k}\,^{pq}\,{}_{l})\nabla_{p}\nabla_{q}f_{G}$ (27)
where $f_{G}=\frac{df}{dG}$. In a FRW space-time it is a perfect fluid tensor. For the spatially flat case, the Gauss-Bonnet invariant is $G=24(\dot{H}H^{2}+H^{4})$ and the Friedmann equations of $f(G)$ gravity are expressible as (eq.5 in [36])
$\kappa\mu=3H^{2}-\kappa(Gf_{G}-f-24H^{3}\dot{G}f_{GG})$ (28)
$\kappa p=-3H^{2}-2\dot{H}+\kappa(Gf_{G}-f)$ (29)
$\qquad-16\kappa H(H+\dot{H})f_{G}-8\kappa H^{2}\ddot{f}_{G}$
With $\dot{f}=f_{G}\dot{G}$ we obtain $\dot{f}_{G}=f_{GG}\dot{G}$ and the first equation rewrites as $\kappa\mu=3H^{2}-\kappa(Gf_{G}-f-24H^{3}\dot{f}_{G})$.
Comparison with (20) gives the following identification
$\mathscr{B}=\kappa(Gf_{G}-f-24H^{3}\dot{f}_{G})$ (30)
After straightforward calculations we infer
$-\mathscr{A}=\frac{\dot{\mathscr{B}}}{H}=8\kappa[\dot{f}_{G}(H^{3}-2H\dot{H})-H^{2}\ddot{f}_{G}].$ (31)
Thus from (21) we obtain eq.(29).
Proposition 3
The Friedmann equations of Cotton gravity with the perfect fluid Codazzi tensor (13) are the Friedmann equations of Gauss-Bonnet $f(G)$ gravity with the choice (30).
III.3 $\mathbf{f(T)}$ gravity
In the framework of gravity theories with torsion, the “teleparallel equivalent of general relativity” is the best known one. It is widely discussed in [5] and briefly reviewed in [24]. It is based on the following action (see [24] eq. 2.5):
$S=\frac{1}{2\kappa}\int d^{4}x\sqrt{-g}\left[T+f(T)\right]+S^{(m)}$
The field equations are eqs.263 in [5] (and eq.2.6 in [24]). For the spatially flat FRW space-time, the Friedmann equations of $f(T)$ gravity are expressible as (eqs 2.9, 2.10 in [24] or eqs.267, 268 in [5]):
$H^{2}=\frac{\kappa}{3}\mu-\frac{f}{6}-2f_{T}H^{2}$ (32)
$\dot{H}=-\frac{1}{2}\frac{\kappa(p+\mu)}{1+f_{T}-12H^{2}f_{TT}}$ (33)
The torsion scalar is $T=-6H^{2}$ (eqs.269 in [5]), and $f_{T}=df/dT$. Comparing with eq.(20) we identify
$\mathscr{B}=-\frac{1}{6}f(T)-2f_{T}(T)H^{2}$ (34)
Note that $\dot{f}(T)=f_{T}\dot{T}=-12H\dot{H}f_{T}$, $\dot{f}_{T}=f_{TT}\dot{T}=-12H\dot{H}f_{TT}$. Thus ˙$\dot{\mathscr{B}}=-2H\dot{H}f_{T}+24H^{3}\dot{H}f_{TT}$ and
$-\mathscr{A}=\frac{\dot{\mathscr{B}}}{H}=-2\dot{H}f_{T}(T)+24H^{2}\dot{H}f_{TT}(T)$ (35)
Summing the Friedmann equations (20) and (21) of Cotton gravity we get
$\frac{\kappa}{2}(p+\mu)=-\dot{H}+\frac{\dot{\mathscr{B}}}{2H}$ (36)
Inserting (35) in (36) gives eq.(33).
Proposition 4
The Friedmann equations of Cotton gravity with the perfect fluid Codazzi tensor (13) are the Friedmann equations of $f(T)$ gravity with the choice (34).
III.4 Einsteinian Cubic and $\mathbf{f(P)}$ gravity
In [4] an extended theory is proposed, based on an invariant $P$ constructed with cubic contractions of the Riemann tensor. The theory was subjected to three constraints: 1) the spectrum should be identical to that of GR (whence the name); 2) it is neither topological nor trivial in d=4; 3) it is independent of the dimension.
The action is
$S=\int d^{4}x\sqrt{-g}\left[\frac{R-2\Lambda}{2\kappa}+P\right]+S^{(m)}$
$P=-\beta_{1}R_{j}\,^{pq}\,{}_{k}R_{p}\,^{rs}\,{}_{q}R^{j}\,_{rs}\,{}^{k}+\beta_{2}R_{jk}\,^{rs}R_{rs}\,^{pq}R_{pq}\,^{jk}$
$+\beta_{3}R^{j}{}_{k}R_{pqrj}R^{pqrk}+\beta_{4}R_{pqrs}R^{pqrs}+\beta_{5}R_{jkpq}R^{kp}R^{jq}$
$+\beta_{6}R_{k}^{p}R_{p}^{j}R_{j}^{k}+\beta_{7}RR_{pq}R^{pq}+\beta_{8}R^{3}$ (37)
The aforementioned constraints impose three linear relations among the coefficients $\beta_{i}$.
In [16] the cosmological applications of Einsteinian cubic gravity at early and late times were investigated. The paper [34] the authors analyze the viability of the theoretical model by considering different observational features, as cosmic chronometers data, baryon acoustic oscillations, and supernovae.
The field equation may be written in the form $R_{kl}-\frac{1}{2}Rg_{kl}+\Lambda g_{kl}=\kappa(T_{kl}+H_{kl})$ where $H_{kl}$ is an involved symmetric tensor containing contractions of the Riemann and the Ricci tensor.
The Friedmann equations are eqs.11 and 12 in [16]:
$3H^{2}=\kappa(\mu+6\alpha\widetilde{\beta}H^{6})+\Lambda$ (38)
$3H^{2}+2\dot{H}=-\kappa\left[p-6\alpha\widetilde{\beta}H^{4}(H^{2}+2\dot{H})\right]+\Lambda$ (39)
with $\widetilde{\beta}=-\beta_{1}+4\beta_{2}+2\beta_{3}+8\beta_{4}$. Comparison with (20) yields
$\mathscr{B}=6\kappa\widetilde{\beta}H^{6}$ (40)
Then
$-\mathscr{A}=\frac{\dot{\mathscr{B}}}{H}=12\kappa\widetilde{\beta}H^{4}\dot{H}$ (41)
From eq.(21) we get eq.(39).
Proposition 5
The Friedmann equations of Cotton gravity with the perfect fluid Codazzi tensor (13) generate the Friedmann equations of cubic Einsteinian gravity with the choice (40).
In the same paper [16] the authors proposed the extension of Einsteinian cubic gravity:
$S=\int d^{4}x\sqrt{-g}\left[\frac{R}{2\kappa}+f(P)\right]+S^{(m)}$
The field equations are still of the type $R_{kl}-\frac{1}{2}Rg_{kl}=\kappa(T_{kl}+\widetilde{H}_{kl})$ where $\widetilde{H}_{kl}$ is quite more involved. The corresponding Friedmann equations are eqs. 26, 27 in [16]:
$3H^{2}=\kappa\mu-\kappa f-18\kappa\widetilde{\beta}H^{4}(H\partial_{t}-H^{2}-\dot{H})f_{P}$ (42)
$3H^{2}+2\dot{H}=-\kappa P-\kappa f-6\kappa\alpha\widetilde{\beta}H^{3}$ (43)
×$\qquad\times\left[H\partial_{t}^{2}+2(H^{2}+2\dot{H})\partial_{t}-3H^{3}-5H\dot{H}\right]f_{P}$
where $f_{P}=\frac{df}{dP}$ and $P=6\widetilde{\beta}H^{4}(H^{2}+2\dot{H})$.
It is simple to identify
$\mathscr{B}=-\kappa f-18\kappa\widetilde{\beta}H^{4}(H\dot{f_{P}}-H^{2}f_{P}-\dot{H}f_{P})$ (44)
Now $\dot{f}(P)=f_{P}\dot{P}=18\widetilde{\beta}H^{3}(4H^{2}\dot{H}+4\dot{H}^{2}+H\ddot{H})f_{P}$. After tedious but straightforward calculations it is
˙$\frac{\dot{\mathscr{B}}}{H}=$ $6\kappa\widetilde{\beta}H^{3}[2H\dot{H}f_{P}-\dot{f}_{P}(4\dot{H}-H^{2})-H\ddot{f}_{P}]$ (45)
Thus from (21) we get (43).
Proposition 6
The Friedmann equations of Cotton gravity with the perfect fluid Codazzi tensor (13) are the Friedmann equations of $f(P)$ gravity with the choice (44).
III.5 Regularized cubic Lovelock gravity in four dimensions
In Section 3 of [7] the authors focused on the four dimensional cubic Lovelock gravity in a FRW space-time. They obtained the following Friedmann equations
$\kappa\mu=3J^{2}(1+\alpha J^{2}+\beta J^{4})$ (46)
$\kappa(p+\mu)=-2(\dot{H}-\frac{R^{\star}}{6a^{2}})(1+2\alpha J^{2}+3\beta J^{4})$ (47)
where $J^{2}=H^{2}+\frac{R^{\star}}{6a^{2}}$. We dropped their cosmological constant and the stress energy tensor is multiplied by a factor 2 to match with our notation. We can state the following:
Lemma 7
If in (21) it is $-\mathscr{B}=F(J^{2})$ where $F$ is a smooth and arbitrary function of $J^{2}=H^{2}+\frac{R^{\star}}{6a^{2}}$ then
$\frac{\kappa}{2}(p+\mu)=-(\dot{H}-\frac{R^{\star}}{6a^{2}})[1+F_{J}(J^{2})]$ (48)
where $F_{J}=\frac{dF}{dJ^{2}}$.
Proof. From (21) we have $\frac{\kappa}{2}(p+\mu)=\frac{R^{\star}}{6a^{2}}-\dot{H}+\frac{\dot{\mathscr{B}}}{2H}$. If $-\mathscr{B}=F(J^{2})$ then ˙$\dot{\mathscr{B}}=-F_{J^{2}}(2J\dot{J})$; on the other hand $J\dot{J}=H(\dot{H}-\frac{R^{\star}}{6a^{2}})$ so that ˙$\frac{\dot{\mathscr{B}}}{2H}=-(\dot{H}-\frac{R^{\star}}{6a^{2}})F_{J^{2}}$ and the Lemma is proven. normal-$\square$
Now take $F_{J}(J^{2})=\alpha J^{2}+\beta J^{4}$ and (47) are recovered.
III.6 Sussman Nájera Model in Cotton gravity
In [43] eq.16, the authors introduced the following modified Friedmann equation:
$H^{2}=\frac{\kappa}{3}\mu-\frac{R^{\star}}{6a^{2}}-\frac{\gamma}{a^{2}}\mathcal{K}(t)$ (49)
with an arbitrary dimension-less function $\mathcal{K}(t)$ and a constant inverse square lenght $\gamma$. Then they computed the components of the Codazzi tensor that solves Cotton gravity. Comparison with (20) gives
$\mathscr{B}=-\frac{\gamma\mathcal{K}(t)}{a^{2}}$ (50)
$\mathscr{A}=-\frac{\dot{\mathscr{B}}}{H}=\frac{\gamma\dot{\mathcal{K}}(t)}{a^{2}H}-\frac{2\gamma\mathcal{K}(t)}{a^{2}}$ (51)
The Codazzi tensor reads explicitly
$\mathscr{C}_{kl}=\left[\frac{\gamma{\dot{\mathcal{K}}}(t)}{a^{2}H}-\frac{2\gamma\mathcal{K}(t)}{a^{2}}\right]u_{k}u_{l}-\frac{\gamma\mathcal{K}(t)}{a^{2}}g_{kl}$ (52)
and compares with the components in eq.17 in [43].
IV Comparison with conformal Killing gravity
After Cotton gravity, Harada introduced a new theory of gravity to explain the present accelerated phase of the Universe without explicit introduction of dark energy [22]:
$\nabla_{[j}R_{kl]}-\frac{1}{3}\nabla_{[j}Rg_{kl]}=\nabla_{[j}T_{kl]}-\frac{1}{6}\nabla_{[j}Tg_{kl]}$ (53)
The equations are manifestly of third order in the derivatives of the metric tensor.
Shortly after, we introduced in [33] a parametrization of the theory by showing that (53) are equivalent to the Einstein’s equations in which the stress-energy tensor is augmented by a divergence-free conformal Killing tensor:
$R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+K_{kl}$ (54)
$\nabla_{[j}K_{kl]}=\frac{1}{6}\nabla_{[j}Kg_{kl]}$ (55)
where $K=g^{pq}K_{pq}$. We named this theory conformal Killing gravity.
The second equation defines a divergence-free conformal Killing tensor (CKT). They are deeply investigated in differential geometry and in physics [14, 26, 39, 42].
We proved existence of a conformal Killing tensor in any FRW space-time, obtaining two modified Friedmann equations that allow for the presence of a dark sector. When applied to a simple toy model, this theory reveals a phantom dark fluid with equation of state (EoS) parameter $w=-5/3$ [33].
In a second paper [23] Harada developed an interesting cosmological analysis confirming in general that the dark energy predicted by the conformal Killing gravity has the same EoS parameter.
In this Section we investigate the connections between Cotton and Conformal Killing gravity.
To this end, consider a generic space-time endowed with a (0,2) symmetric tensor satisfying the relation
$\nabla_{j}K_{kl}=a_{j}g_{kl}-b_{k}g_{jl}-b_{l}g_{jk}$ (56)
We call such tensors Sinyukov-like (see [29],[42]). If $K_{kl}$ is the Ricci tensor we recover the Sinyukov manifolds, investigated for example in [17].
Here we consider divergence-free Sinyukov-like tensors:
$\nabla_{j}K_{kl}=5\frac{\nabla_{j}K}{18}g_{kl}-\frac{\nabla_{k}K}{18}g_{jl}-\frac{\nabla_{l}K}{18}g_{jk}$ (57)
where $K=K^{p}{}_{p}$. They satisfy the condition (55) that defines divergence-free conformal Killing tensors (CKT) [39].
A space-time with a Sinyukov-like divergence-free tensor is a solution of conformal Killing gravity (54).
On the other hand (57) implies the Codazzi condition
$\nabla_{j}\left[K_{kl}-\frac{K}{3}g_{kl}\right]=\nabla_{k}\left[K_{jl}-\frac{K}{3}g_{jl}\right]$ (58)
So $\mathscr{C}_{kl}=K_{kl}-\frac{1}{3}Kg_{kl}$ is a Codazzi tensor with $\mathscr{C}_{r}^{r}=-\frac{1}{3}K$. From $R_{kl}-\frac{1}{2}Rg_{kl}=T_{kl}+K_{kl}$ we recover the paradigm (7) with the same stress-energy tensor.
Proposition 8
A space-time with a divergence-free Sinyukov-like tensor (57) is a solution both of conformal Killing gravity (54) and of Cotton gravity (7), with the same stress-energy tensor.
Rather surprisingly, we show that any FRW space-time is equipped with a Sinyukov-like tensor.
We recall that a vector $Z_{j}$ is a Conformal Killing vector [39] (CKV for short) if the following condition holds:
$\nabla_{j}Z_{i}+\nabla_{i}Z_{j}=2\psi g_{ij}$ (59)
where the scalar function $\psi$ is called conformal factor.
Let $Z_{j}=Fu_{j}$ with $u_{j}u^{j}=-1$ and $F$ a scalar function. The following result holds:
In a GRW space-time $Fu_{j}$ is a CKV if and only if $\dot{F}=HF=\psi$ and $\nabla_{i}F=-u_{i}\dot{F}$, i.e. $F$ depends only on time. ([15] thrm 2.1, [35] thrm 1)
In this case, since $H=\dot{a}/a$, we obtain $F(t)=ka(t)$ for some constant $k$. According to Rani et al. [39], the CKV originates a conformal Killing tensor
$K_{ij}=F^{2}u_{i}u_{j}+F_{1}g_{ij}$ (60)
for arbitrary scalar function $F_{1}$. Let us choose $F_{1}$ in order that $0=\nabla_{p}K^{p}{}_{j}$. A simple evaluation using (9) shows that $\nabla_{i}F_{1}=-5F\dot{F}u_{i}$. Then $F_{1}$ depends only on time, and $\dot{F}_{1}=5F\dot{F}$. An integration gives $F_{1}=\frac{5}{2}F^{2}-\Lambda$ being $\Lambda$ a constant. Now
$K_{jk}=F^{2}(u_{j}u_{k}+\frac{5}{2}g_{jk})-\Lambda g_{jk}$ (61)
The evaluation of the gradient gives: $\nabla_{i}K_{jk}=HF^{2}(-5u_{i}g_{jk}+g_{ij}u_{k}+g_{ik}u_{j})$. Contraction with $g^{jk}$: $\nabla_{i}K=-18HF^{2}u_{i}$. It turns out that $K_{jk}$ satisfies (57), i.e. it is divergence-free Sinyukov-like. $\square$
The associated Codazzi tensor $\mathscr{C}_{kl}=K_{kl}-\frac{K}{3}g_{kl}$ is
$\mathscr{C}_{ij}=F^{2}(u_{i}u_{j}-\frac{1}{2}g_{ij})+\frac{\Lambda}{3}g_{ij}$ (62)
We have proven
Proposition 9
Any GRW space time, and thus any FRW space-time, is a solution of both Cotton and conformal Killing gravity with the same stress-energy tensor.
The Codazzi tensor (62) is not as general as (13), since it is fixed up to a constant. In fact the condition (14) is more general than $\dot{F}=HF$. In a FRW space-time Cotton gravity is more general than conformal Killing gravity.
The CKT (61) is used in [33] to obtain the Friedmann equations of conformal Killing gravity. The eigenvalue equation $K_{ij}u^{i}=\lambda u_{j}$ gives $\lambda=\frac{3}{2}F^{2}-\Lambda$ and thus $K=6\lambda+2\Lambda$. We rewrite the tensor as
$K_{ij}=\frac{2\lambda+2\Lambda}{3}u_{i}u_{j}+\frac{5\lambda+2\Lambda}{3}g_{ij}$ (63)
Note that $2\lambda=3F^{2}-2\Lambda=3k^{2}a^{2}(t)-2\Lambda$.
V Fixing the dark sector
From the above discussion it is clear that the dark sector is described by the Codazzi terms and emerges from geometry. In Cotton gravity the term $\mathscr{B}$ remains unfixed, so that further restrictions are needed.
We make a standard cosmological analysis by supposing that the content of energy in $T_{jk}$ is from radiation (r) and matter (m): $\mu=\mu_{r}+\mu_{m}$ where
$\mu_{r}=\frac{\mu_{r,0}}{(a/a_{0})^{4}},\quad\mu_{m}=\frac{\mu_{m,0}}{(a/a_{0})^{3}}$
By setting $\frac{8\pi G}{3H_{0}^{2}}=1/\mu_{c}$, $\Omega_{r,0}=\mu_{r,0}/\mu_{c}$, $\Omega_{m,0}=\mu_{m,0}/\mu_{c}$, $\Omega_{k,0}=-\frac{R^{\star}}{6H_{0}^{2}a_{0}^{2}}$ and $\Omega_{\Lambda}=\frac{\Lambda}{3H_{0}^{2}}$ we get
$\frac{H^{2}}{H_{0}^{2}}=\frac{\Omega_{r,0}}{(a/a_{0})^{4}}+\frac{\Omega_{m,0}}{(a/a_{0})^{3}}+\frac{\Omega_{k,0}}{(a/a_{0})^{2}}+\frac{\mathscr{B}}{H_{0}^{2}}+\Omega_{\Lambda}$
In terms of red-shift $1+z=a_{0}/a$ the equation becomes
$\frac{H^{2}}{H_{0}^{2}}=\frac{\Omega_{r,0}}{(1+z)^{4}}+\frac{\Omega_{m,0}}{(1+z)^{3}}+\frac{\Omega_{k,0}}{(1+z)^{2}}+\frac{\mathscr{B}(z)}{H_{0}^{2}}+\Omega_{\Lambda}$ (64)
If $\mathscr{B}=0$ the standard $\Lambda$CDM model is recovered.
It is quite remarkable that we only need to suppose the presence of matter and radiation, while the theory provides the term that can be interpreted as a dark sector. As Harada argued the dark sector appears as a purely geometric effect due to the presence of the Codazzi tensor. In more detail, let’s write the condition ˙$\dot{\mathscr{B}}=-H\mathscr{A}$ as a function of the redshift.
With $\dot{z}=-(1+z)H$ it is ˙$\dot{\mathscr{B}}=\frac{d\mathscr{B}}{dz}\dot{z}=-\frac{d\mathscr{B}}{dz}(1+z)H$. The condition becomes
$\mathscr{A}=\frac{d\mathscr{B}}{dz}(1+z)$ (65)
In this representation $\mathscr{A}$ does not depend on the Hubble parameter.
Now recall eqs.(18) and (19): $\kappa\mu_{D}=3\mathscr{B}$ and $\kappa p_{D}=-3\mathscr{B}-\dot{\mathscr{B}}/H$. Suppose that an EoS $p_{D}=w(z)\mu_{D}$ is valid, where the parameter $w$ may be redshift-dependent.
In general the dark sector is characterized by $w<-1/3$. The regime $-1<w<-1/3$ is usually called “quintessence”, while the one with $w<-1$ is called “phantom”. The consequences of a phantom energy in the Universe were pointed out in the seminal paper [6].
The EoS and (65) imply the equation
$3\mathscr{B}(1+w(z))=(1+z)\frac{d\mathscr{B}}{dz}$
with solution
[…]
Inserting this in (64) we have
$\frac{H^{2}}{H_{0}^{2}}=\frac{\Omega_{r,0}}{(1+z)^{4}}+\frac{\Omega_{m,0}}{(1+z)^{3}}+\frac{\Omega_{k,0}}{(1+z)^{2}}+\Omega_{\Lambda}$ (67)
[…]
This is substantially eq.14 in [2] with the difference that here $\Lambda$ is not dynamical. We also note the balance
$1=\Omega_{r,0}+\Omega_{m,0}+\Omega_{k,0}+\Omega_{\Lambda}+\Omega_{D,0}$ (68)
where $\mathscr{B}_{0}/H_{0}^{2}=\Omega_{D,0}$ is the present-time dark energy density.
This analysis generalizes the considerations in [23, 33]. In particular, if $w(z)=w$ we get the wCDM model with cosmological constant reviewed in [2]:
$\mathscr{B}(z)=\mathscr{B}_{0}(1+z)^{3(1+w)},\quad\mathscr{A}(z)=3(1+w)\mathscr{B}(z)$
Reversing to cosmic time we get
$\mathscr{B}(t)=\frac{\mathscr{B}_{0}}{(a(t)/a_{0})^{3(1+w)}}$ (69)
$\mathscr{A}(t)=3(1+w)\mathscr{B}(t)$ (70)
In the case $w=-5/3$ we recover the phantom term typical of conformal Killing gravity discovered in [22, 23, 33].
The Codazzi tensor becomes
$\mathscr{C}_{kl}=\frac{\mathscr{B}_{0}}{(a/a_{0})^{3(1+w)}}\left[3(1+w)u_{k}u_{l}+g_{kl}\right]$ (71)
There are many redshift-dependent models that parametrize the shape of dark energy: they were used for example in [2] to test deviations from the $\Lambda$CDM model. More recently they were discussed on the base of JWST results [28]. The same parametrizations can be used to fix $\mathscr{B}$ and $\mathscr{A}$ using (66). We recall some of them here.
V.1 Chevallier-Polarski-Linder (CPL) model
It is one of the most used redshift-dependent parametrisation, and was introduced in [12, 27]. It supposes that
$w(z)=w_{0}+w_{a}\frac{z}{1+z}$
where $w_{0}$ is the present time dark energy EoS and the correction describes its evolution. It features a good behavior at high $z$ and it is linear at low $z$ (see[1] and [28] for details). From (66) we obtain
$\mathscr{B}(z)=\mathscr{B}_{0}(1+z)^{3(1+w_{0}+w_{a})}\exp(-\frac{3w_{a}z}{1+z})$ (72)
$\mathscr{A}(z)=3\mathscr{B}(z)(1+w_{0}+\frac{w_{a}z}{1+z})$ (73)
The CPL model has a counterpart in the Codazzi parametrization of Cotton gravity.
Remark 10
In [1] the authors observe that the recent data from JWST reveal a very large number of massive galaxies at high redshift. This fact poses challenges to the standard $\Lambda$CDM model. Based on the CPL model and testing with the new datasets, they propose a scenario in which the Dark Sector consists of a negative cosmological constant.
V.2 Jassal-Bagla-Padmanabhan (JPB) model
In [25] the authors introduced the following expression for the EoS parameter
$w(z)=w_{0}+w_{a}\frac{z}{(1+z)^{2}}$
They claim that this form solves some issues present in the CPL model (see [28] and references therein). From (66) again we easily find
$\mathscr{B}(z)=\mathscr{B}_{0}(1+z)^{3(1+w_{0})}\exp\left[\frac{3}{2}\frac{w_{a}z^{2}}{(1+z)^{2}}\right]$ (74)
$\mathscr{A}(z)=3\mathscr{B}(z)\left[1+w_{0}+\frac{w_{a}z}{(1+z)^{2}}\right]$ (75)
Also this model has a counterpart in the Codazzi parametrization of Cotton gravity, without explicit introduction of dark energy.
VI Conclusions
Cotton gravity offers a simple setting to reproduce the Friedmann equations of well known extended theories. In all cases the dark sector arising from geometry is described by a Codazzi tensor with the proper choice of a single function. We also showed that the recently proposed conformal Killing gravity is absorbed in Cotton gravity at least for cosmological FRW space-times. The dark sector may be fixed requesting an EoS: this can accomodate in a unified description the best known redshift dependent models.
Data availability
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.